Suppose S = 1dk +-+--+. . . + (1+k)n+1 k-0.07, and n-20. Round your final answer to two decimal places. btain the value of S if d-4, g-0.01,
Answer #1

S = 4/(1+0.07) + 4*(1+0.01)/(1+0.07)2 + 4*1.012/1.073......+ 4*1.0120/1.0721 =

S = 4/1.07 *(1 + 1.01/1.07 + (1.01/1.07)2.....(1.01/1.07)20

S= 4/1.07 * ((1.01/1.07)20-1)/(1.01/1.07-1) = 46.82

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well